OpenStudy (anonymous):

solve usng multiplication principle -7/8x = -9/10

OpenStudy (anonymous):

What's the multiplication principal?

OpenStudy (anonymous):

-7/8x = -9/10 -7*10x=-9*8 -70x=-72 x=72/70 x=36/35

OpenStudy (anonymous):

I'm just waiting for someone to tell me to "google it"...